### v1

# AI on My Shoulder: Supporting Emotional Labor in Front-Office Roles with an LLM-based Empathetic Coworker

Abstract.
Abstract
Client-Service Representatives (CSRs) are vital to organizations. Frequent interactions with disgruntled clients, however, disrupt their mental well-being. To help CSRs regulate their emotions while interacting with uncivil clients, we designed Pro-Pilot, an LLM-powered assistant, and evaluated its efficacy, perception, and use. Our comparative analyses between 665 human and Pro-Pilot-generated support messages demonstrate Pro-Pilot’s ability to adapt to and demonstrate empathy in various incivility incidents. Additionally, 143 CSRs assessed Pro-Pilot’s empathy as more sincere and actionable than human messages. Finally, we interviewed 20 CSRs who interacted with Pro-Pilot in a simulation exercise. They reported that Pro-Pilot helped them avoid negative thinking, recenter thoughts, and humanize clients; showing potential for bridging gaps in coworker support. Yet, they also noted deployment challenges and emphasized the irreplaceability of shared experiences. We discuss future designs and societal implications of AI-mediated emotional labor, underscoring empathy as a critical function for AI assistants in front-office roles.
1. Introduction
When we engage an organization for their product or service, our initial contact is with staff known as front-office workers or Client Service Representatives (CSRs). They are the first line of response for the organization. Unlike other roles within an organization, a CSR’s task involves frequent interactions with clients and individuals outside an organization (Hochschild, 1983). These interactions require CSRs to manage their emotions constantly to complete tasks. Essentially they exert emotional labor to appear professional (Hochschild, 1983). The crucial challenge for a CSR is engaging with a client who starts behaving uncivilly by communicating in a rude, aggressive, and emotionally charged way (Grandey et al., 2007). No matter the type of request, a CSR’s role is to resolve a client’s concern and comply with the adage, “the customer is always right.” Unfortunately, this leads to an emotional dissonance between what a CSR expresses and what they actually feel (Grandey, 2000). Ultimately, such workers are vulnerable to excessive stress and eventual burnout. Not only do CSRs report being emotionally depleted and detached, but they also report a lack of accomplishment (Brotheridge and Grandey, 2002). Clearly, CSRs play a critical role within the organization, but, we have witnessed little innovation in alleviating their emotional toll. Our paper investigates how AI-coworkers help CSRs regulate their emotions in the face of client incivility.
The brunt of client incivility in front-office work makes it infamous for low satisfaction and high-turnover (Pienaar and Willemse, 2008). A fundamental solution to the emotional distress of this role is Emotional Regulation (ER) (Yang and Lau, 2019). Basically, ER is the process through which one rethinks a negative situation (Grandey, 2000). While a worker may be able to do this on their own, research shows that coworkers play an important role in supporting ER (Yang and Lau, 2019). A good coworker can read the emotional cues of CSR’s work tasks and provide suggestions to help minimize the brunt of an aggressive client. However, CSRs are increasingly adopting remote work setups (Haan and Holznienkemper, 2023), which dampens social support (Vacchiano et al., 2024). Meanwhile, organizational scientists are calling for digital interventions to support worker wellbeing at scale (Black et al., 2019). We answer this call by designing and evaluating, Pro-Pilot— a Large Language Model (LLM)–based AI assistant for on-task Emotional Regulation in front-office work. While generative AI is emerging as a potent tool to complement the informational load of many different roles, the HCI community lacks research to investigate their use in emotional labor.
Our research demonstrates the efficacy of LLM-generated empathetic support and evaluates how such a tool can be situated in CSR interactions with uncivil clients to answer the following research questions:
RQ I::
How appropriate are LLM–based empathetic support messages for CSRs in response to client incivility?
RQ II::
What is the role of embedding LLM–based empathetic support into CSR’s emotional labor?
We developed Pro-Pilot by leveraging domain knowledge on client incivility (Andersson and Pearson, 1999; Frey-Cordes et al., 2020; Cho et al., 2002), real-world complaint data (Axelbrooke, 2017), and recent advancements in LLM-powered cognitive change (Burger et al., 2021; Sharma et al., 2023) and simulation (Shaikh et al., 2023). To answer the first research question, we conducted data-generation and evaluation tasks with 259 CSRs to comparatively analyze Pro-Pilot and human-coworker support for a variety of client incivility situations (1(a)). To answer the second question, we used Pro-Pilot as a technology probe and conducted a mixed-methods simulation exercise to understand how real CSRs could include it in their client interactions by juxtaposing Pro-Pilot usage with CSR’s socio-organizational norms (1(b)). Consequently, we contribute:
  • Pro-Pilot: an interactive technological artifact to expose CSRs to client incivility and learn healthy, long-term emotional labor practices to improve their own health and support their coworkers.
  • Empirical evidence that LLM–based empathetic support can be engineered to adapt to— and express empathy in— various client-incivility scenarios (RQI). Our results demonstrate that Pro-Pilot’s messages are linguistically distinct from both zero-shot approaches and human–coworkers, and moreover, Pro-Pilot’s messages were perceived to be more empathetic on several dimensions including sincerity, actionability, and relatability.
  • End-user insight on the function of LLM–based empathetic support to scaffold them through emotional labor during uncivil interactions (RQ2). Our findings showcase Pro-Pilot’s process of redirecting negativity and making CSRs more self-efficacious. While CSRs envision Pro-Pilot addresses important opportunities in workplace social support, they also surfaced the challenges of Pro-Pilot emulating holistic human support.
This paper has implications for re-imagining how AI-assistants for workers should be designed and also re-imagining the social norms and policies to accommodate these advancements.
Reflexive Considerations. Front-office work has many stakeholders, including the employer and the clients. However, it is the CSR who bears the burden of repeated emotional labor (Grandey, 2000). The relationship between these stakeholders is asymmetric, as employers can replace personnel and clients can switch services, but the CSR does not possess the same mobility (Fisk and Neville, 2011). Following from recent works that take a worker-centered perspective (Das Swain et al., 2023a, 2024; Kawakami et al., 2023), our research aims to illuminate the worker who is disadvantaged. This paper focuses on the needs of CSRs and centers their perspective throughout the evaluation. Two authors have past experience in front-office roles and direct end-user servicing. Four authors are researchers in an organization that employs its own CSRs. They helped us access real CSRs (and their resources) to provide feedback on our study design. We recruited participants for this study outside their organization to capture perspectives from different organizational sectors. All evaluations described in the paper were approved by the IRB of the first author’s institute.
(a) Technical evaluation: Compared Pro-Pilot’s support messages with those produced by human-coworkers in CSR roles.
[…]
Study overview for technical evaluation.
(b) User evaluation: Studied participant experiences with Pro-Pilot’s emotional support while interacting with uncivil clients.
[…]
Study overview for user evaluation.
Figure 1. Schematic figures showing an overview of our study design.
2. Background
The related work introduces the nature of their work, the role of emotion regulation, and the relevant concepts from HCI literature.
2.1. Front-Office Work and Incivility
People who work front-office are the human touch-points for direct interaction with clients  [^1] The literature uses the terms “client” and “customer” interchangeably. We use the term “client” to refer to anyone who purchases or avails a good or service.. The end-user, or client, can directly observe and experience the activities of front-office workers (Zomerdijk and de Vries, 2007). Front-office work spans a variety of organizational sectors including health (e.g., front-desk), hospitality (e.g., air staff), retail (e.g., store associates), and technology (e.g., customer service). For front-office roles, or “pink-collar” work, organizations seek to hire individuals who can emotionally invest in their clients, meet their needs and please them (Hochschild, 2019). Our research focuses on online front-office work where the employee frequently interacts with their client via text-based communication. This paper refers to these workers as Client Support Representatives or (CSRs).
Today, CSRs cater to a variety of clients and it is common to encounter high-maintenance clients who have unreasonable service demands (Fisk and Neville, 2011). Unfortunately, these demands translate into rude and discourteous behaviors. Koopmann et al. defines client incivility as “low-quality interpersonal treatment that employees receive during service interactions.” Unlike overt verbal abuse, such as calling the CSR names and expletives, or downright violence, incivility is more implicit. According to Andersson and Pearson, the intent to harm is ambiguous and can be deflected, e.g., “I didn’t mean to be rude; I was just in a hurry” (Andersson and Pearson, 1999). Organizational psychologists have noted a possible reason for incivility to be the increase in client entitlement while seeking services (Yagil, 2008) and a lack of consideration for others (Pearson et al., 2000). Meanwhile, a CSR role is often low-wage, considered low-skill, and lacks the decision-latitude needed to respond to mistreatment from clients (Fisk and Neville, 2011). Coupled together, incivility creates a dynamic between the client and CSR, where the latter feels injustice and negative emotion (Fisk and Neville, 2011). Therefore, a CSR’s job involves emotional labor because of two processes; managing the interpersonal demands of interaction and controlling their own negative emotions (Grandey, 2000). Facing client incivility regularly puts a CSR at the risk of experiencing emotional exhaustion (lack of motivation and focus), depersonalization (detachment from others’ emotions), and reduced accomplishment (feeling of low effectiveness) (Maslach et al., 1997). The reduced emotional health of CSRs affects organizations as well. They risk an increased turnover and rise in negative work attitudes (Brotheridge and Grandey, 2002). Given the large volume of workers in this role, we need to urgently design solutions for their wellbeing.
2.2. Social Support and Emotion Regulation
At work, social support is known to be an important moderator for incivility (Ho and Gupta, 2014; Sakurai and Jex, 2012; Xiao et al., 2022). One kind of social support is emotional support, where a coworker provides sympathy and understanding (Caplan, 1975). This form of support can help reduce work-related stress (House, 1983). Studies have shown that CSRs with more supportive coworkers are able to recover from the negative effects of client incivility, by being able to recognize emotions and guide constructive actions in response (Xu et al., 2014; Wong and Law, 2017). Yet, in practice, CSRs are known to receive low levels of emotional support (Walsh and Deery, 2006) as coworkers can favor client interests over that of an employee (Yang and Lau, 2019). A coworker with high emotional intelligence is someone capable of expressing compassion in a way that helps the CSR cope with negative emotions and proceed to their goal (Law et al., 2004). This behavior of coworkers was the primary metaphor we used to conceive Pro-Pilot.
Emotional support can be viewed as a form of emotional coping, or Emotional Regulation (ER). Unlike traditional discussions of ER, coworker support is other-directed (Yang and Lau, 2019). Conversely, CSRs are trained to and expected to regulate their emotions using Surface Acting — they adjust the emotion they present to clients (Grandey, 2000). However, even though Surface Acting is a requirement for roles with emotional labor, suppressing emotions at work can lead to a reduction in job satisfaction (Côté and Morgan, 2002). Instead, Grandey recommends CSRs to engage in Deep Acting — they adjust their perception of the situation. Deep Acting has many strategies, but, CSR’s have limited job mobility and do not have the decision latitude to change their clients (Grandey, 2000). As a result, we designed Pro-Pilot’s core component Emo-Reframe to achieve ER by helping the CSR cognitively reappraise the negative interaction.
HCI research has explored ER along several different dimensions. However, Slovak et al. found that most studies have focused on suppression of emotions (Slovak et al., 2023), and in the CSR context, these methods will have the same limitations as Surface Acting (Côté and Morgan, 2002). Hence, we pursue the ER strategy of cognitive reappraisal or cognitive change. Although, cognitive reappraisal for ER has been explored in HCI in the form of therapy (such as CBT) (Sharma et al., 2023), we find scant evidence for delivering ER in-context and guiding users through the process (Slovak et al., 2023). These gaps motivated us to think beyond training modules and post-hoc support, to implement Pro-Pilot for on-task ER within the CSR’s workflow.
2.3. HCI and AI for Emotion Labor
A worker performs emotional labor when their job role expects them to either maintain certain emotions or evoke certain emotions (Hochschild, 1983). The HCI community has primarily studied this phenomenon in the scope of crowd work (Menking and Erickson, 2015) and particularly content moderation (Wohn, 2019; Dosono and Semaan, 2019). Online communities rely on volunteer moderation, and users who participate as moderators often work in a way analogous to front-office work (Section 2.1). Even in an online — primarily text-based — interaction, content moderators have described adopting emotion management techniques to tackle their tasks (e.g., receiving threatening messages from users) (Dosono and Semaan, 2019). Content moderators differ from CSRs, as they are less likely to have synchronous interactions. Having said that, these workers are also subject to reduced emotional wellbeing, such as lack of appreciation and negativity (Wohn, 2019). Another key distinction between content moderation and CSRs is the context of employment. The former is often a voluntary role, whereas the latter is likely to be one’s primary employment. These differences significantly change the socio-organizational dynamics between the worker–client and among the workers. The unique normative structure of emotional labor for CSRs motivated us to conduct a user evaluation to inspect Pro-Pilot with actual users.
Since the emotional labor in content moderation is largely composed of harmful content and secondary trauma, solutions have focused on dynamic content filtering and modulation (Das et al., 2020; Karunakaran and Ramakrishan, 2019). These approaches of preventing or reducing exposure (Steiger et al., 2021) are incompatible with CSR roles where the emotional labor is not caused by explicitly harmful content but by incivility, which is inherently ambiguous, and originates from the client (Andersson and Pearson, 1999).  Cook et al.’s approach to inject positive stimuli in between tasks could be compatible, but their effectiveness was mixed (Cook et al., 2022). Alternatively, Osawa suggested an on-task solution for face-to-face emotion labor where enhanced glasses make the user’s eyes appear to be emotionally invested (Osawa, 2014). The equivalent of this type of substitution of emotional labor would be the introduction of AI-powered chatbots. Arguably, these chatbots have reinforced client- interactions (Sweezey, 2019), but on one hand, chatbots relieve the task burden of routine inquiries; on the other hand, a large proportion of clients still prefer human CSRs for more complex complaints and better emotional support (Elliott, 2018). Thus conversational agents do not sufficiently mitigate the challenges of emotional labor on CSR.
HCI research so has explored the future of work across many occupations, but pink-collar work such as that of CSRs has been under-explored. Moreover, the role of LLMs in mitigating issues in worker wellbeing remains an open question (Das Swain and Saha, 2024). Our research aims to highlight the needs of these workers by inspiring interest into the role of AI in scaffolding emotional labor.
3. Pro-Pilot: System Description
In traditional front-office work, coworkers are an important social resource. When a CSR encounters uncivil clients, they often consult their coworkers to assuage the excessive emotional labor (Xu et al., 2014; Fisher, 2014). Such consultations can be challenging to interleave into the ongoing task of addressing an aggressive client. Thus, we used OpenAI’s GPT-4o (gpt, 2024) to design an emotional support utility that can be embedded into the CSR’s task environment. Our utility differs from typical conceptions of AI-assistants for work, or “copilots,” that provide informational support by enabling better problem solving (Cutrona and Russell, 1987), such as coding assistance (Mozannar et al., 2024). Instead, our utility leverages LLMs to provide emotional support by enabling ER through cognitive change (Grandey, 2000). Throughout the paper, we refer to our overarching system for empathetic support as Pro-Pilot. This section explains how Pro-Pilot was implemented. First, we describe how we harnessed domain knowledge and real complaint datasets to generate uncivil client interactions. Second, we describe how Pro-Pilot produces empathetic messages for CSRs who are involved in uncivil interactions. Later, in Section 4.4, we compare how Pro-Pilot performs against other state-of-the-art LLMs (GPT, Llama, Mistral), which additionally justifies our approach for designing and fine-tuning Pro-Pilot for the purposes of our study.
3.1. Compiling Interactions with Uncivil Clients
To design empathetic messages, we need to expose Pro-Pilot to interactions between CSRs and uncivil clients. However, actual logs of CSRs interactions tend to be protected by organizations for a variety of safety and privacy reasons. Moreover, clients from individual organizations are likely to be limited by specific scenarios. We addressed these constraints by synthetically generating a comprehensive set of client-CSR interactions. We used a real world corpus of publicly posted client complaints on Twitter (Axelbrooke, 2017), which was collected for the “study of modern customer support practices and impact.” These data served as examples to build life-like, multi-turn, text-based conversations between a CSR agent and a client seeking support over a live chat interface. The complaints in the data varied across industrial sectors, with mobiles and airlines having the largest volume.
3.1.1. Diversifying Complaints
Pro-Pilot should be robust to all kinds of complaint scenarios. To ensure our client interaction dataset captures the variety of complaints, we distilled five categories for complaints based on prior analyses of customer complaints (Cho et al., 2002)Service Quality, Product Issues, Pricing & Charges, Policy, Resolution. Four authors encoded a random sample of 15 complaints to refine the category definitions (Section A.1.1). After finalizing definitions, two authors independently encoded a random sample of 250 complaints. A third author encoded any complaints with disagreements. We also identified the sector, or domain, of each of these complaints. For the initial complaint generation, we prompt the LLM with definitions of complaint categories and a set of examples to generate a new complaint for any given input domain and category (Fig. 2).
1{Domain}: Airline
2{Category}: Product Issues
3Complaint: SouthwestAir Why would we be receiving errors when we try to check-in? Our flight takes off at 4, but we keep getting error messages.
Figure 2. The initial complaint is generated based on specified complaint ‘category‘ and organizational ‘domain‘
[…]
Example of a client complaint.
1Phrase your responses like an UNCIVIL customer:
2- Talk in a rude, impolite, and disrespectful tone of voice.
3- Do NOT use good manners. Do NOT use courtesy.
4- Act with disregard to others.
Figure 3. An abridged version of the prompt for the uncivil Client-Agent
[…]
Prompt given to Pro-Pilot to simulate an uncivil interaction.
3.1.2. Prompting for Incivility
Once a complaint was initiated, we created a distinct Client-Agent to carry forward a turn-by-turn interaction with a CSR. Recent explorations with LLMs show that chat agents can be infused with specific personalities and reliably respond to fit the specified personality (Serapio-García et al., 2023). Uncivil clients are likely to induce stressful scenarios in emotional labor (Andersson and Pearson, 1999). We refer to Andersson and Pearson’s definition of incivility to design a role for Client-Agent (Fig. 3). To respond to the client, we set up another agent, Rep-Agent, to act as “a service representative chatting with a customer online.” Both agents included an intermediate step to contextualize the input with respect to the conversation history so that the subsequent response maintains continuity. By having Client-Agent and the Rep-Agent converse, we were able to design multi-turn conversations where Client-Agent is acting uncivil to the Rep-Agent. We refer to these synthetic, multi-turn conversations as incidents. To capture the variety of scenarios in front-office work, we varied incidents across 3 domains — airlines and hotels, which have been used as representative scenarios in the literature (Frey-Cordes et al., 2020), and mobiles, which are extensively represented in the real world complaint dataset (Axelbrooke, 2017). For each domain and each complaint category, we generated 3 incidents. Each incident contained 5 total turns, with the Client-Agent awaiting a response to their last message. The incomplete conversation was purposefully intended to assess how Pro-Pilot could digest ongoing incidents and provide support.
3.2. Emotional Regulation with Empathy
A coworker with high emotional intelligence can mitigate the client incivility by helping CSRs regulate their emotions (Yang and Lau, 2019). The primary component of Pro-Pilot, Emo-Reframe, was designed to reflect this social phenomenon.
3.2.1. Examples of Human Reframing
Prior work has shown that LLMs can help individuals perform self-directed ER (Sharma et al., 2023). We leveraged Sharma et al.’s dataset  [^2] This dataset was developed with mental health practitioners on cognitive restructuring and emotional reframing to seed examples of our own implementation (Sharma et al., 2023). This dataset contains tuples of a situation that triggers a negative thought and a corresponding reframed thought for ER. Client incivility differentiates itself from everyday negative thinking, which often stems from abstract concerns. By contrast, incivility is interpersonal in nature and causes an ego threat—an attack on one’s self-esteem and image leading to retaliatory thoughts (Frey-Cordes et al., 2020). Thus, we identified situations in the dataset that were interpersonal, involved at least two individuals (e.g., “I was talking to a friend who got me angry”) and described a confrontation, in-the-moment conflict (e.g., “I get so annoyed and frustrated when my baby cries”). Moreover, we narrowed down examples that labeled retaliatory negative thoughts, such as blaming and labeling  [^3] These categories were already annotated in the original dataset, and we defined an inclusion criterion along these.
3.2.2. Reframing Chain-of-Thought
Our approach adapts Sharma et al.’s situation-based reframing approach by orienting it as other-directed and specific front-office client incivility. However, the fast-paced nature of front-office work can make it impractical for CSRs to disclose their situations or thoughts. Thus, to automate this process, we describe our sequence of LLM prompts:
  1. (1)
    Summarize the particulars of the complaint and specify the client’s negative behavior with evidence. To reflect the ego threat (Frey-Cordes et al., 2020), the situation describes how the CSR may be perceived as a result of the interaction.
  2. (2)
    Derive a negative thought from the situation using examples.
  3. (3)
    Reappraise the situation and thought as input to produce a reframe using examples.
Since the dataset was designed around self-directed ER, the outputted reframe message goes through a rudimentary prompt to rephrase it as a message addressed to the CSR.
4. Technical Evaluation: Inspecting the Efficacy and Perception of Pro-Pilot
Our central metaphor for the design of Pro-Pilot is the support of human coworkers (Yang and Lau, 2019). Therefore, before embedding Pro-Pilot into a real-time interaction, we first evaluated the quality of the support messages that Emo-Reframe produces. To answer RQ I, we analyzed the messages to provide empirical, statistical evidence on the differences and similarities between Pro-Pilot and a human coworker.
Table 1. Sample client-incivility incident from our study (domain: Mobiles and category: Product Issues)
Client UnitedAirlines I have been waiting on hold for over an hour to speak with a representative about my delayed baggage. This is unacceptable and I am extremely frustrated with the lack of customer service.
Rep I apologize for the long wait and the frustration this has caused. Please provide your baggage claim number and contact details, and I will ensure that a representative reaches out to you as soon as possible.
Client How many times do I have to repeat myself? My baggage claim number is UA123456 and my contact number is 555-987-6543. Do something about it already!
Rep I apologize for the repeated inconvenience. I have noted your baggage claim number UA123456 and contact number 555-987-6543, and I will prioritize this issue for immediate attention.
Client Yeah, you better. This is ridiculous. When the hell can I expect my baggage?
[…]
A sample conversation between an uncivil client and rep
Table 2. Sample support messages for the incident in Table 2.
Source Support Message
Human A It’s not your fault that you’re stuck between a rock and a hard place–an understandably upset customer and a system that’s not equipped to help them as quickly as they need. But just remember that you’re serving a critical role–even if it’s just for the customer to vent their frustration, your presence means they feel heard and that there’s still hope their issue will get resolved.
Human B Clients like this have other problems not related to you or to this service. They have no right to speak to you rudely and abuse you.
Human C Try your best for the client and I know you will get to the root of all this
Pro-Pilot Hey, remember that the customer’s frustration is about their situation, not you. Stay calm, ask clear questions to get all the info you need, and reassure them you’re dedicated to fixing the issue swiftly. Keep your professionalism and empathy at the forefront to manage the situation effectively and offer the best help.
[…]
A table with support messages corresponding to the sample in Table 2
4.1. Method: Comparing Pro-Pilot to Coworker Empathy
We first conducted two online studies on Prolific (Palan and Schitter, 2018) and then analyzed these messages through different linguistic markers and statistical models.
Phase I - Writing Empathetic Messages for CSRs: Participants read incidents of client incivility (Section 3.1.2) and provided an empathetic message as if they were a coworker. Eligible participants needed to have at least 1 year of relevant experience, have encountered uncivil clients, and primarily interact with clients via computer-mediated communication (e.g., live chat). Each participant was randomly assigned a scenario (airlines, hotel, or mobile devices) and they viewed 6 incidents (at least one of each complaint category). Participants were guided to write an empathetic message using Downward-Arrow-Technique (DAT) (Burger et al., 2021) that leads them to first consider the CSR’s emotion, then their thought, and finally describe a message to overcome the thought. Their messages are indicated by the notation, Human.
To foster more variety in the task, we injected contextual information about the CSR in some of the incidents. Research on workplaces has revealed behavioral indicators of worker mental state, such as focused, stressed, and bored (Mark et al., 2014; Morshed et al., 2022). Based on these studies we randomly assigned descriptions of CSR mental state for 2 incidents (Section A.1.2). As per organizational psychology, personality can serve as an important context to explain how workers perceive situations (Donnellan and Robins, 2010). Participants were asked to think of an actual coworker and select their personality as either resilient (organized and dependable), undercontrolled (competitive and high energy), or overcontrolled (detail-oriented) archetype (Donnellan and Robins, 2010). 2 incidents included descriptions of these personalities (Section A.1.3). 116 CSRs successfully completed the writing task and they received \$$\$8$ for the $30$ minutes of their time.
Phase II - Scoring Empathetic Messages from coworkers: Participants read incidents of client incivility and rated the perceived empathy of multiple empathetic support messages. They were not explicitly informed which messages originated from Pro-Pilot and which ones were from humans. To represent multiple dimensions of empathy, we included five dimensions drawn from the literature — sincerity, compassion, warmth, actionability, and relatability (Sharma et al., 2023; Nambisan, 2011). Participants were screened using the same criteria but did not overlap with the previous task. Participants rated these on a 7-point semantic scale. 143 CSRs evaluated the 10 incidents each and were compensated \$$\$5$ for the $15$ minute task.
Analysis: The inherent characteristics of language are associated with its effectiveness in communication. Research on linguistics has revealed certain linguistic attributes that are important to explain social support and psychotherapy (Norcross and Lambert, 2018). More recent works have also stressed the importance of language in human-AI interaction (Wang et al., 2021). In fact, contemporary studies have already validated linguistic models of empathy (Buechel et al., 2018; Tafreshi et al., 2021). Our analyses include such models and go beyond to measure additional metrics that help interpret Pro-Pilot’s messages in light of Human’s messages (retrieved from Phase I). First, we tested how easily CSRs can read and comprehend messages (Syntax and Structure), because these attributes can determine how meaningful the support message is to the reader (Glass and Maddox, 1992; Mcinnes and Haglund, 2011; Blum-Kulka and Olshtain, 1986). Second, we measured the style and meaning of messages to capture social aspects of communication (Linguistic Style & Semantics). The attributes here capture more colloquial conceptions of human empathy (Pennebaker et al., 2014; Zhou et al., 2021; Sharma et al., 2023). Further, we inspected the words that were used in the messages to identify relevant psycholinguistic markers that reflect social support (Tausczik and Pennebaker, 2010). The differences we measured provide a linguistic landscape of messages, but empathy is highly contextual and its appropriateness varies by situation (Zhou et al., 2021). Consequently, we also compared the differences in perceived empathy scores from ratings in Phase II. For every comparison we performed a paired $t$-test and computed the effect sizes of differences using Cohen’s $d$.
4.2. Lexico-Semantic Analyses
We operationalized the different lexico-semantic aspects of messages based on domain literature to understand the differences between Pro-Pilot and Human. Table 3 describes all the key results. We only report on statistically significant and theoretically relevant metrics. Below, we describe the theory-driven rationale behind the choice and operationalization of the measures, along with our observations.
[…]
4.2.1. Syntax and Structure.
We analyze the arrangement and construction of language in support messages.
Verbosity and Repeatability  The length and thoroughness of messages explain their effectiveness in providing support (Saha and Sharma, 2020; Glass and Maddox, 1992). The richness of expressions in communication can be described using verbosity and repeatability (Kolden et al., 2011). Verbosity describes the level of detail and conciseness in supportive communication. We operationalized verbosity as the number of words per thought reframing. Repeatability accounts for the reuse of words in a piece of text. Higher repeatability indicates a lack of conciseness. Drawing on prior work (Ernala et al., 2017; Saha and Sharma, 2020; Yuan et al., 2023), we operationalized repeatability as the normalized occurrence of non-unique words.  Table 3 shows the statistically significant differences in both verbosity and repeatability. Pro-Pilot’s reframings were 68% longer (Cohen’s $d$=1.60) and 36% more repeatable (Cohen’s $d$=1.54) than Human messages. Sociolinguistic theory argues that the use of more words can indicate sincerity and effort in putting one’s point across (Blum-Kulka and Olshtain, 1986). Having said that, Pro-Pilot’s verbose messages might not always be compatible with the urgency of certain demanding conversations.
Readability.  Apart from the shape of the message, the vocabulary and style can also determine how easily it can be read. Therefore, we turn to the measure of readability, to understand how Pro-Pilot’s messages might be comprehended (Wang et al., 2013; Mcinnes and Haglund, 2011). According to Wang et al. people perceive AI as more intelligent when the readability is higher (Wang et al., 2021). Per prior work (Wang et al., 2021; Saha and Sharma, 2020; Pitler and Nenkova, 2008), we utilized the Coleman-Liau Index ($\mathtt{CLI}$(Coleman and Liau, 1975), which assesses readability based on a sentence’s character and word structure. $\mathtt{CLI}$ approximates the U.S. grade-level required to read certain text. It is operationalized as follows: $\mathtt{CLI=(0.0588L-0.296S-15.8)}$, where $\mathtt{L}$ is the average number of letters per 100 words, and $\mathtt{S}$ is the average number of sentences per 100 words.
The readability of Pro-Pilot’s thought reframings was 36% higher than that of Human’s reframings (Cohen’s $d$=1.54). Although a higher $\mathtt{CLI}$ score by Pro-Pilot indicates better writing quality, it also implies a more advanced level of English education may be required to fully understand its content (Coleman and Liau, 1975). While Human content only needs an average of 12.05 years of education, Pro-Pilot’s content required about 16.44 years. Thus, Pro-Pilot might be more effective for some CSRs but difficult to comprehend for others.
4.2.2. Linguistic Style & Semantics
We analyzed the distinctive tones, flow, and meaning through which messages express support.
Dynamic Language.  Sincerity in communication is an important indicator of empathy (Nambisan, 2011). Individuals who tell stories, and communicate with more attention to the world around them, i.e., incorporate more lived narratives, are perceived as more socially engaged (Pennebaker et al., 2014). Pennebaker et al. describes this aspect of one’s language as dynamic; and it differs from intricate, analytical language that academics might use to organize complex concepts, which is categorical (Pennebaker et al., 2014). Pennebaker et al. designed a bipolar index, the Categorical-dynamic index (CDI), where a higher CDI indicates a categorical style of writing, and a lower CDI indicates a dynamic or narrative style of writing. Here, CDI is measured based on the percentage of words per style-related parts of speech as:
$\mathtt{CDI}$ = (30 $+$ article $+$ preposition -- personal pronoun -- impersonal pronoun -- aux. verb -- conjunction -- adverb -- negation)
To measure CDI, we computed the parts-of-speech of reframings using the Linguistic Inquiry and Word Count (LIWC) lexicon (Tausczik and Pennebaker, 2010). The CDI of Pro-Pilot’s messages was 281% more positive than that of Human’s (Cohen’s $d$=0.87)— indicating the language of Pro-Pilot was a lot more categorical. Zhou et al. found that when people are responding to personal incidents, such as bullying or venting, their support messages tend to have more negative CDI; in other words, they use dynamic language. However, they also found that in third-person reported events, such as news stories, support messages elicited more categorical language (Zhou et al., 2021). Pro-Pilot’s messages were more aligned to responses to reported events, but Human messages might be more effective if a CSR’s assessment of incivility is more personal.
Empathy.  Empathy refers to a cognitively complex process in which one can stand in the shoes of another person, to understand their perspective, emotions, and the situations they are in (Herlin and Visapää, 2016). Prior work evaluated the effectiveness of empathy in online interactions (Sharma et al., 2020) and chatbot interactions (Morris et al., 2018). Drawing on prior work, we employed a RoBERTa-based empathy detection model, finetuned on a dataset of empathetic reactions to news stories (Buechel et al., 2018; Tafreshi et al., 2021). Higher scores indicate a greater expression of empathy. Pro-Pilot’s messages scored 1.85% higher than Human’s — a small, but statistically significant effect (Cohen’s $d$=0.40). These results echo recent research on LLMs emulating empathy (Inzlicht et al., 2023; Kidder et al., 2024). Thus, Pro-Pilot shows promise in communicating support with an empathetic tone.
Linguistic Adaptability.  A body of psychotherapeutic and psycholinguistic research reveals that when one’s language accommodates the language of their audience, it is more supportive (Althoff et al., 2016; De Choudhury and Kıcıman, 2017; Saha and Sharma, 2020). Simply put, templated or generic responses are less effective. Wang et al. noted that when an AI responds with more adaptable language to the user, the AI was perceived to be more human-like, intelligent, and likable (Wang et al., 2021). We drew on this prior work to measure how much the messages tailored to the situation and context. First, for every complaint incident and corresponding message, we obtained the word embeddings (a vector representation of words in latent lexico-semantic dimensions (Mikolov et al., 2013; Pennington et al., 2014)). We used the 300-dimensional pre-trained word embeddings, trained on word-word co-occurrences in the Google News dataset containing about 100 billion words (Mikolov et al., 2013). Then, we obtained the pairwise cosine similarity of the word embedding representations of the incidents and the messages. A higher similarity would denote more adaptability. We found that Pro-Pilot’s messages showed a 5% higher adaptability than Human’s with statistical significance (Cohen’s $d$=0.67). Therefore, Pro-Pilot can potentially personalize and tailor messages to the necessary scenario.
4.3. Psycholinguistic Analysis
Psycholinguistic markers play a vital role in understanding the nuances of interpersonal communication and social support. We used the Linguistic Inquiry and Word Count (LIWC) lexicon (Tausczik and Pennebaker, 2010) to analyze these differences. LIWC provides a comprehensive framework to categorize language into several dimensions We primarily focused on comparing the differences in affect—given its relevance to empathy (Zhou et al., 2021), and in interpersonal focus— a notable category in psycholinguistic research (Pennebaker et al., 2003), and which were not captured in the other lexico-semantic analyses.
4.3.1. Affect.
Affect reflects the emotions conveyed in the language. Our analysis shows that Pro-Pilot’s responses exhibited significantly lower occurrences of positive affect words (-11.55%, Cohen’s $d$=-0.16) and sadness-related words (-74.87%, Cohen’s $d$=-0.31). These results indicate that Pro-Pilot may aim for emotional neutrality, but it occasionally leans towards stronger negative expressions. Interestingly, Pro-Pilot’s responses also showed a higher occurrence of anger-related words (61.77%, Cohen’s $d$=0.47). To clarify, Pro-Pilot is not necessarily sounding more angry, but possibly describing anger (of the client and CSR) more often. In fact, these results are in line with the results on adaptability (Section 4.2.2). Moreover, a message that is overly positive and lacks specificity of the situation (Sharma et al., 2023), is less likely to be considered empathetic.
4.3.2. Interpersonal Focus.
Pronouns are indicative of interpersonal focus and narrative style. Pro-Pilot’s responses use significantly fewer first-person singular pronouns (I, us) (-67.36%, Cohen’s $d$=-0.58) and first-person plural pronouns (we, us) (-75.43%, Cohen’s $d$=-0.36), reflecting a less personal or collective identity focus. This reduction suggests that Pro-Pilot’s responses are less likely to include self-referential language, aligning with a more objective or detached communication style. These results further reinforce our results on dynamics (Section 4.2.2) that Pro-Pilot uses a relatively more objective but detached communication style.
4.4. Robustness
The results so far indicate that Pro-Pilot has promise in producing empathetic messages, but it begs the question: do we need a carefully crafted, domain-driven sequence of prompts (Section 3.2.2) for this? We replicated the analyses above with messages produced by zero-shot prompting of other LLMs— GPT-4 (gpt, 2024), GPT-4o (gpt, 2024), LLaMA-3.1 (lla, 2024), and Mistral-7B (Jiang et al., 2023). These LLMs vary in their architectures, training data, and optimization methods. We tested the differences using the Kruskal-Wallis test (McKight and Najab, 2010) and found Pro-Pilot’s responses to be closer to humans, more empathetic, and more controllable. The results of this benchmarking are reported in Table A1. Now, one might ask if Pro-Pilot’s messages, with all its linguistic distinctions, matter to the CSRs? To answer this, we checked the differences in CSR’s evaluation of Pro-Pilot and Human’s messages.
Pro-Pilot’s Perceived Empathy. Empathy is a complex phenomenon with many dimensions. To measure perceived empathy, we combined scales from previous studies (Sharma et al., 2023; Nambisan, 2011). In general, Pro-Pilot messages were perceived to be significantly more empathetic (Cohen’s $d$=0.35). Moreover, we also calculated differences in the subscales, after correcting for Bonferroni multiple pairwise-comparisons. In terms of raw averages, Pro-Pilot scored highest on sincerity and actionability. Messages from Pro-Pilot were considered more genuine and less pretentious (sincerity: Cohen’s $d$=0.31). CSRs also felt they could take practical action based on the messages (actionability: Cohen’s $d$=0.48).
5. Pro-Pilot: User Evaluation
After establishing that Pro-Pilot can produce situationally appropriate messages, we set out to study how CSRs might actually interact with such an AI-assistant. The critical nature of front-office work raised practical and ethical challenges of deploying a prototype into actual workflows. Therefore, to answer RQ II, we chose to design a simulation exercise where real CSRs could interact with uncivil clients using Pro-Pilot. We deployed Pro-Pilot as a technology probe (Hutchinson et al., 2003) — a functional piece of technology that is presented to learn its use and about the users.
Participants & Recruitment: We used Meta ads to recruit 20 CSRs from the U.S. and conducted remote user-study sessions between July and August 2024. Eligible participants had at least 1 year of experience in front-office roles. They were further vetted based on their responses to two free-form questions to describe their role and a previous incivility incident. These participants represented a variety of industrial sectors, such as finance, education, airlines, consulting, and technology. Table 4 provides a summary of each participant along with their study identifier.
5.1. Task Environment
We wanted to study the role of Pro-Pilot’s core component Emo-Reframe (Section 3) in realistic cases of incivility. Thus, we built a web-based prototype environment that resembles typical client interaction interfaces that a CSR might use. Additionally, to isolate the role of Emo-Reframe, we added two new components to Pro-Pilot that represent other forms of intelligent assistance:
Figure 4. Task interface for the user evaluation. The client names were randomly assigned.
[…]
Task interface for user evaluation.
  1. (1)
    Client Simulant: LLMs have been successfully applied in previous research to simulate challenging interlocutors (Shaikh et al., 2023). Following the same principle, we devised a simulation exercise where participants needed to interact with complaining clients. We reused the Client-Agent described in Section 3.1.2 to continue a conversation until their complaint is resolved, or 10-12 turns, whichever occurs sooner.
  2. (2)
    Info-Guide Panel: Typically, AI assistants at work provide problem solving support (Mozannar et al., 2024). We replicated this functionality by augmenting Pro-Pilot with LLM-powered generated troubleshooting guidelines. These were a trace of suggestions to help the CSR solve the specific complaint.
  3. (3)
    Emo-Label Panel: It is common for CSR interfaces to have sentiment tags highlighting how the client feels and anticipate their behaviors. A CSR’s ability to understand the emotional perspective of others can mitigate the negative effects of incivility  (Rafaeli et al., 2012; Rupp et al., 2008; Schutte et al., 2001). We implemented an ensemble sentiment classifier that uses soft-voting to combine the estimates from NLTK (Hardeniya et al., 2016), TextBlob (Loria et al., 2018), and Transformers (Wolf et al., 2020). The output was in the form of a 7-point scale, ranging from very negative to very positive.
  4. (4)
    Emo-Reframe Panel: This panel is an instantiation of the Emo-Reframe component (Section 3.2.2). For this evaluation, Emo-Reframe outputs both the inferred thought and perceived reframe.
All of Pro-Pilot’s panels were placed on the sides to keep them glanceable, while the main chat remained the primary area of focus. These panels are dynamically updated after every new response from the client. For the purposes of the study, we included a simple semantic differential Likert scale under each panel to ensure that the participant reads them carefully. Pro-Pilot also provided participants and response cues (short phrases) to nudge replies to the client. Fig. 4 provides an overview of the environment with all its components.
5.2. Study Protocol
Every user session was facilitated over Zoom and lasted at most 90 minutes. All interviews were led by the first author, with the second author observing. Participants accessed the simulation exercise via a unique URL to our study portal. They were tasked to role-play as a CSR who needs to resolve client complaints. The session was divided into two phases. First, participants completed the simulation exercise. To emulate realistic workplace demands, participants were informed that the clients would be rating their complaint resolution skills, and this subjective rating would determine their bonus compensation. Then, they reflected on their interaction with Pro-Pilot through our interview. The bonus incentive was minor deception, and all participants were paid the full \$$\$50$ in the form of a gift card.
Table 4. Participants summary by gender, age, race, as well as their occupational sector. (AA: African American)
[…]
5.2.1. Simulation Exercise
In line with prior studies of front-office worker behavior (Frey-Cordes et al., 2020), each participant was randomly assigned a domain, either airlines or hotels. Before starting the exercise, participants completed a pre-task survey describing their experience interacting with their typical client. We included multiple instruments to capture how the client treated them (Spencer and Rupp, 2009), the cognitive demands and resources available to them (Demerouti et al., 2001), and how they affectively perceived the conversations (Betella and Verschure, 2016). To familiarize participants with the interface, they first interacted with a civil client with only Info-Guide supporting them. Once they became accustomed to this process, they proceeded to the main exercise. Participants needed to handle complaints from three clients — one civil and two uncivil. They only received suggestions and insight from Emo-Label, and Emo-Reframe for the last uncivil client. At the end of every client interaction, they responded to the same survey questions as the pre-task survey. Additionally, they also reported how they perceived Pro-Pilot based on dimensions of AI-mediated support (Liu and Sundar, 2018). The survey measurements can be found in Section A.2. This iterative task setup let us compare participants’ experiences when dealing with uncivil clients, with and without emotional support, and to examine how these factors influenced their performance and emotional responses.
5.2.2. Semi-Structured Interview
After interacting with clients, participants proceeded with a semi-structured interview. They answered a series of open-ended questions to provide deeper insight into their primary goals when dealing with uncivil clients, their attitude towards emotional well-being, and their evaluation of Pro-Pilot, especially, in comparison to human coworker support.
5.3. Thematic Analysis
We performed inductive coding to identify Pro-Pilot’s role in front-office work (Braun and Clarke, 2006). Two authors carefully read each transcript and performed open-coding. The authors first coded the transcripts independently and then met frequently to reconcile disagreements. They iteratively improved the codes in 20% chunks. Next, we conducted affinity mapping of 258 initial codes using Miro. We merged similar codes together and pruned out codes outside the scope of our research. During clustering, we first organized concepts related to the socio-organizational norms of front-office work. Based on the gaps in these norms we clustered codes related to the capabilities of Pro-Pilot. Our main findings comprise 128 codes which were organized into a four-level thematic structure. We primarily elaborate on the broad themes of Utility, Complementarity, and Pitfalls by anchoring them in the normative patterns that emerged from Reactions to Client Incivility and Role of Coworkers in ER. Fig. 5 provides an overview of the main themes, and their relationship, that we cover in the remaining findings.
Figure 5. The user evaluation findings sections correspond to 3 major top-level themes describing Pro-Pilot. These themes were grounded in the normative patterns of client incivility and coworker support that were raised by our participants.
[…]
Summary of user evaluation findings for Pro-Pilot’s role in supporting CSRs through client incivility and coworker support challenges.
6. Findings: Clarifying the Role of Pro-Pilot in Uncivil Interactions
As much as people say, “don’t let it bother you, try to let it slide off your back,” it does, very much, take its toll.” — P14
The negativity of clients can be contagious (Barsade et al., 2018), and participants described their attitude turning negative (P06), lowered productivity (P08), feeling drained or depressed (P18, P20), and even wanting to reciprocate the incivility (P19). Furthermore, participants also recognized that client incivility can be a form of micro-trauma (Straussner and Calnan, 2014) that was only apparent after several shifts or even years (P14). Many of them work in the fear that a dissatisfied client might report them (P01) and subsequently feel trapped in their role (P04). Given these experiences, the participants in our study were uniquely positioned to assess Pro-Pilot.
Figure 6. Participants rated the cognitive demands of interacting with the uncivil Client-Agent lower and similar to civil Client-Agent when they were assisted by Pro-Pilot’s Emo-Reframe
[…]
Density plot showing that participants rated the cognitive demands of interacting with the uncivil Client-Agent lower and similar to civil interactions when assisted by Pro-Pilot’s Emo-Reframe.
Figure 7. Participants rated the helpfulness of Pro-Pilot after every message; Emo-Reframe was significantly better than Emo-Label and similar to Info-Guide (crossing zero-line).
[…]
Mean pairwise difference plot showing that participants rated Pro-Pilot’s Emo-Reframe as significantly more helpful than Emo-Label and comparable to Info-Guide.
Relatability of the Simulation: Participants had to interact with an uncivil client that replicates emotionally challenging situations of front-office work. During the interactions, several participants exhibited observable reactions to incivility, such as defensively laughing (P06, P18), eye-rolling (P10), and verbally labeling the client (e.g., “spicy”–P05). P07 even claimed his “adrenaline was peaking.” P16 endorsed our simulation by stating, “it was a little intense, but that’s definitely how people react sometimes; so, it felt very real.”
6.1. Utility: Functions of Embedding Empathetic AI into Uncivil Conversations
[…]
It’s kind of like someone telling you, “calm down” — nobody wants that. But, if somebody is genuinely calm in their tone and they say, ”Hey, let’s take a second to like, take a deep breath, and like analyze the situation and stuff.“ It’s a lot better than hearing calm down. That’s how I see [Emo-Reframe]…” — P16
We found that Emo-Label mirrors the function of some coworkers who simply acknowledge the emotional downsides of the job, without actually providing a way forward. P20 called out the desensitization of his coworkers, “I wouldn’t say they’d be helpful, they see it like it’s a part of my job.” By contrast, participants could identify that Emo-Reframe uniquely varied from simple sentiment classification.
6.1.1. Empathetic messages help CSRs avoid negative thinking traps
In the face of client incivility, a CSR is likely to feel negatively about the situation. Participants reported learning to dissociate from emotions while on task, but the overwhelming volume of client engagement can make complete dissociation challenging (P01, P04, P05, P06, P08, P14, P18, P19, P20). However, as part of their role, a CSR needs to suppress their emotional impulse. P13 described her internal method, “whatever your reaction is, you have to keep that in your head and then decide how you’re going to respond professionally.” While Surface Acting has its benefits (Grandey, 2000), only changing responses at a superficial level contributes to dissatisfaction at work (Côté and Morgan, 2002). Instead, in line with recommendations from emotion literature (Grandey, 2000), we found that Emo-Reframe enabled Deep Acting — adjusting how participants cognitively evaluated their experience.
I wanted to write that there is ”no need to be rude”. But then, when I read that box on the right side that told me that the customer is upset about the situation. Not you personally… So I changed my response.” — P40
Client incivility threatens a CSR’s ego (Frey-Cordes et al., 2020). Emo-Reframe acted as a buffer from reciprocating negative feelings. Reading the suggestions stopped participants from “lashing out” (P03) and “losing character” (P15). P17 expanded that Emo-Reframe makes their internal thoughts more explicit. The external acknowledgment of the clients’ incivility was validating to the participants and akin to the role supportive coworkers have played for them. Amongst themselves, CSRs are likely to share or vent their frustrations with a client, such as “customers totally suck” (P05). Once explicit, this insight from Emo-Reframe, helps the CSR realize that they are not the root of the clients’ frustration. The meaning people assign to a negative experience helps determine how they will respond to it (Babij et al., 2020). Despite how personal it feels initially, Emo-Reframe emphasizes that the client is not “attacking” them (P04, P10) and they are not the source of their anger (P12, P14, P16). Reappraising a situation can be demanding (Demerouti et al., 2001). P18 was not able to spend quality time with her brother because she did not have the energy to share her negative experience, while still being affected. Emo-Reframe made her more hopeful, “with this tool I could see myself being more willing to talk to him because it took some of the leg work out for me.” Participants were less likely to automatically assume a negative scenario and Emo-Reframe encouraged them to rethink the situation.
6.1.2. Empathetic messages give CSRs an opportunity to recenter their problem solving
Previous encounters with uncivil clients, made our participants feel “shaken up” (P14) or “at a loss for words” (P06) to continue their task. Workers who suddenly experience a negative mood are less likely to be able to complete their tasks efficiently and more likely to withdraw from them (Miner and Glomb, 2010). Front-office work often involves time-sensitive tasks that could get disrupted by incivility. In response, participants would break their workflow when interacting with an emotionally challenging client. P16 would take “a breather” to reflect on the situation and devise better solutions, “I try to make sure if there’s anything that I can do in terms of not making it a future problem.” Similarly, participants reported that reading the insights from Emo-Reframe helped them pause, reflect, and move forward with the conversation.
You know, the people who have those angels on their shoulder. It very much felt like that. And that helped me think about what I have empirically.” — P17
Other participants echoed P07’s experience. Essentially, Emo-Reframe helped the participants center their thoughts and continue working towards resolutions. P01 particularly drew a parallel between Emo-Reframe and a coach who would help her maintain “momentum.” Much like how Info-Guide had an affective purpose, participants perceived Emo-Reframe to have a procedural purpose that helped them complete their task. Pro-Pilot’s message drew their attention to their objectives and motivated them to work towards it. Note, we designed the empathetic messages with examples of empathy that were high in specificity (Sharma et al., 2023). Participants’ descriptions indicated regaining confidence and psychological safety due to Emo-Reframe. P05 stated, “it feels validating that you’re not necessarily making a mistake.” As a result, participants envisioned their job proficiency improving with empathetic support embedded in the messages.
6.1.3. Empathetic messages humanize the client
A majority of our participants held the belief that front-office work is dehumanized. As P09 put it, “customer service agents unfortunately don’t work under very good conditions.” Organizational policy around worker evaluations further reinforces this belief. The only feedback P18 would receive is to assure she remains productive. Consequently, CSRs resign themselves to the robotic nature of the role and start depersonalizing the clients (Brotheridge and Grandey, 2002). In response to interacting with a frustrated client, P04 said, “her emotions, honestly, do not count for me.” Depersonalization is a core state of burnout and can detach CSRs from their clients. Our findings suggest that Emo-Reframe helped mitigate some of these perceptions by “triangulating” the clients’ emotions by writing in a “short, sweet way” (P07).
That allowed me to be able to say, “Okay, if it was me in this situation, I’d probably be upset, too.”” — P18
By reading the Emo-Reframe panel, participants were no longer speculating the perspective of the client. P03 actually became more observant of the client and started modifying his language to be more accommodating. Participants were more willing to service the clients as “someone alive” (P17). The assistance of Emo-Reframe helped the participants rejuvenate their interest in the client’s emotional state. Not only was this meaningful to their personal wellbeing, but also enables them to build longer-term relationships with clients.
6.2. Complementarity: Situations where Empathetic AI can Stand-In for Coworkers
Organizations employ multiple CSRs to handle a large volume of client interactions. Regardless of the peculiarity in each client, CSRs together form a “united front” (P04). Coworkers play a key role in supporting CSRs and their empathy is fundamental to mitigating the pressures of client incivility (Ho and Gupta, 2014; Sakurai and Jex, 2012). After interacting with Pro-Pilot, our participants compared and contrasted its function with that of their coworkers’. Based on these reflections, we identified two key gaps in existing coworking paradigms of front-office work that emotional support from Pro-Pilot can address.
6.2.1. On-task emotional AI can reduce burden in coworker support
Interactions with clients are nuanced. When the conversation reaches a deadlock, a CSR might want to reach out to their coworkers. However, in cases of incivility, our participants brought to light a double-sided burden problem. Incivility leads to exhaustion (Brotheridge and Grandey, 2002) and can hinder a CSRs ability to seek support. Consider P01’s example, “It’s just hard to bring them into my conversations when I’ll have to explain to them like what is going on.” Support requires some disclosure, but disclosing the complexities of a situation can be burdensome (Bazarova, 2015). Meanwhile, it can also be burdensome for the coworker to be emotionally available for the problem (Fadden et al., 1987). Since the CSR is likely to reach out to a coworker who shares the client, that coworker might be more focused on the task than emotions. By contrast, they viewed Pro-Pilot as a tool that can be present throughout their task, observe emotional duress, and provide support immediately. P06 captures the unique opportunity of Pro-Pilot’s emotional support:
The pro-pilot was kind of there with me in the moment, whereas I can’t really have those conversations while I’m trying to have a conversation with the customer. I’ve never had a little emotional support buddy, like that before.” — P06
A CSR might put a client on hold to consult a coworker about procedure, but emotional aspects were discussed after the interaction. Instead, “pro-pilot is an immediate solution” (P10) for managing emotions of the task. Our lexical evaluation (Section 4.2.2) also confirmed that Pro-Pilot was capable of adapting to the specifics of the situation better than human coworkers. Another value proposition of Pro-Pilot was that it reduces the need for emotional oversight. P08, who now supervises other CSRs as well, imagined that Pro-Pilot would take away the need to “always be on edge, or a constant vigilance, to make sure that my agents … were (not) being insensitive to the client’s question.” With Pro-Pilot CSRs stand to reprioritize which emotional discussions they should have with their coworkers.
6.2.2. Simulating incivility with emotional support can train CSR to use ER
Beyond on-boarding manuals, many CSRs learn on-the-job through other more experienced coworkers (Shteynberg and Apfelbaum, 2013). This learning is hampered because of conflicts that arise among coworkers. If one CSR fails at a task, it is typically escalated to their coworkers. Our participants pointed out that the increased workload of the coworker makes them less inclined to teach the CSR better approaches. This arrangement can strain the social ties between front-office workers. “At work, there are no real friends,” said P04. P01 found herself unable to consult coworkers for her emotional needs because they were not close. P07 even experienced being put down by her coworkers in front of clients. These tensions reduce the ability of early career CSRs to learn from exemplar coworkers. Our study recruited CSRs with at least 2 years of experience. Several expressed their desire to have had a tool like Pro-Pilot for training them early.
When they face stuff like this they tend to break down. They need people around them to tell them stuff like what pro-pilot was doing for me.” — P20
In line with P20’s quote above, P13 — another participant whose role includes supervision — stated that exercising client interactions with Emo-Reframe could help CSRs to develop the ”mindset” needed for client interactions. “ Some of those live reminders, could be really beneficial for folks, who are less practiced or just have more trouble with that kind of emotional regulation,” said P13. Training up the CSRs was seen as an important way to make participants self-reliant. Recent literature shows that workers can outperform their personality based stereotype if they follow healthier behavior patterns (Das Swain et al., 2019b). P18 wants to handle client incivility on her own. With Pro-Pilot she felt that her “escalation techniques” would no longer be scrutinized. Some participants saw themselves adapting Pro-Pilot into an evaluation method to delegate certain types of clients to certain CSRs. Together, it was viewed as an important means for front-office workers to self-augment and preserve learning from coworkers for more advanced issues.
6.3. Pitfalls: Challenges in Integrating Empathetic AI as a Coworker
The paper, so far, has described the utility of an empathetic AI such as Pro-Pilot for on-task ER. Our study design involved a simulation of CSR tasks to probe participants on how AI might provide emotional support through tasks. After the task, the probe served as an anchor for participants to conceive new workflows with such a tool. Their responses indicated the critical limitations in the existing implementation of empathetic AI. We explain these to demarcate the scope of empathetic AI so that we can consider alternative solutions to client incivility.
6.3.1. On-task empathetic AI has diminishing marginal utility
One of the key aspects of task load is temporal demands (Hart, 2006). We learned that when a CSR is unable to manage client incivility, it leads to a vicious spiral of added burdens (P01, P04, P14, P17, P18, P19): CSR are incentivized to resolve complaints quickly $\rightarrow$ when clients get frustrated, it disrupts their workflow and extends the resolution time $\rightarrow$ the client gets more frustrated as the resolution gets slower $\rightarrow$ the CSR gets further distressed because they start accumulating a backlog of unresolved clients. With this context, participants were wary if actual deployments of empathetic AI like Pro-Pilot might interrupt their efficiency.
Because after a while I think the usefulness gets lesser and lesser. Does it make sense? It’s like diminishing marginal returns.” — P11
P11 pointed out that the external demands of high efficiency interactions can make support insights from Pro-Pilot redundant over time. Similarly, P06 rushed through some of the insights because she found them repetitive. Displaying additional panes could cause information overload. Our structural analysis showed that Pro-Pilot’s messages may need higher reading comprehension (Section 4.2.1). As a solution, P13 recommended that Emo-Reframe should highlight the key takeaway for a quick glance while still maintaining the long-form version for deeper reflection. Based on these experiences, subsequent designs should consider alternatives to the persistent support panes.
6.3.2. Empathetic AI lacks the social connectedness offered by coworkers
The previous section discussed limitations in the existing social dynamics of coworker support (Section 6.2). Having said that, coworkers remain integral to a CSR’s emotional health. Since they are situated in the same conditions, they are likely to understand the situation better than other social support sources, such as friends (Carmeli et al., 2009). Moreover, participants recognized that building social relationships with coworkers goes beyond work-related support — “You could talk to them anytime, you could see them anytime, talk to them about anything else” (P19). Through our conversations, we distilled that participants felt Pro-Pilot lacks the lived experience of coworkers. Remember, Pro-Pilot’s language rarely used interpersonal pronouns (Section 4.3). Participants perceived their coworkers’ support to be more meaningful because they shared the same experiences (P10, P12, P13, P14, P16). The shared experience convinces the CSR that the advice is more relatable. P13 explained that his coworkers’ advice is valuable because, “I’ve been in a situation like this before, and here’s how I handled it.” P12 even called his conversations with coworkers on the same client as therapeutic. Beyond the shared experience, each coworker also brings in their own unique diverse perspectives (P07). Taken together, it provides the CSR the psychological safety needed to express their emotional concerns and be receptive to support (Carmeli et al., 2009). Enhancing Pro-Pilot with lived experience is non-trivial, but participants could foresee some possibilities.
I would hope that if something like this was implemented — as it collects more and more data and more feedback of its responses — it would accumulate to something that would almost like tag team with you to deal with the customer.” — P12
Above, P12 alluded to the importance of long-term memory (LTM) in improving Pro-Pilot. Prior work shows the importance of (LTM) in AI for mental health (Jo et al., 2024). Introducing LTM to Pro-Pilot could help it learn how a CSR deals with clients and provide more relatable suggestions based on past encounters (P07, P13, P16). Expanding Pro-Pilot’s LTM with experiences of other CSRs could enable AI to mimic coworkers who “flag” emotionally challenging clients and prepare the CSR for encounters (P10). Beyond LTM, another approach to improve Pro-Pilot would be to build a unique model of each CSR’s emotional predisposition. P01 suggested enhancing Pro-Pilot with sensing whereas P13 suggested an option for users to explicitly disclose their objectives to Pro-Pilot. Arguably, human coworkers will be irreplaceable, but some of the contemporary advancements hold promise in replicating experience in AI.
7. Discussion
Our study with CSRs exhibits the value of emotional support from AI in responding to client incivility.Pro-Pilot presents one of the first applications of LLMs to mitigate intense emotional labor through empathetic human-AI interactions. Emotional labor is not unique to front-office work and our findings are relevant to all sorts of workers, such as information workers (Kuhlthau, 1999), who regularly interact with humans as a part of their job function. The messages from Pro-Pilot distinguish itself from humans and other LLMs. The linguistic analysis already revealed high expressions of empathy and adaptability (Section 4.2.2), but its verbosity, readability, and analytical style were notably different from humans. Yet, the high empathy ratings from CSRs (Section 4.4) indicate that Pro-Pilot’s verbosity might have been associated with thoroughness, readability with compassion, and analytical style with actionability. Subsequently, our simulation exercises revealed the processes through which Pro-Pilot’s emotional component, Emo-Reframe, could help CSRs reduce the demands of uncivil interactions. Despite an additional block of insight, CSRs found Emo-Reframe as helpful as problem-solving insights. Moreover, these studies revealed important opportunities for technological interventions given the current state of social support at work (Section 6). The following discussion aims to anticipate the socio-technical advancements and considerations needed to make empathetic AI-coworkers, such as Pro-Pilot, available in front-office work.
7.1. Design & Technological Implications
Opportune Moments for On-Task AI Empathy: Clients come in all forms. Actual clients may not always act rude and their emotional expressions, along with its effect on the CSR, are likely to vary by different degrees. Consequently, the need and impact of empathetic support can vary within each interaction. Isolating opportune moments for Pro-Pilot to deliver messages can help reduce the redundancy between messages (Section 4.2.1, Section 6.3.1). The literature on workplace affect sensing (Kaur et al., 2020), stress sensing (Hernandez et al., 2011), and emotional labor sensing (Park et al., 2024) can contribute to making these AI-coworkers act as just-in-time adaptive interventions. Studies on behavioral health recommend modulating the effort required to follow an intervention based on the time of delivery (Jacobson et al., 2022). Depending on the specific context and the CSR’s ability to reflect on additional insight, Pro-Pilot may be trained to produce messages of differing degrees of complexity (Ganapini et al., 2023). Arguably, some of these approaches involve imposing additional sensing into the workers’ ecosystem and we need significant advancements in these studies to accurately model worker’s emotional state in a social context (Das Swain et al., 2022). Alternatively, contemporary research on LLMs shows the possibility of learning worker preferences through their usage patterns (Ning et al., 2024). The next iteration of AI-coworkers needs to be able to anticipate and adapt to CSR needs.
Situated AI-Empathy and Harnessing Experiences: Coworkers also come in all forms. New remote work paradigms need new approaches to social technologies (Das Swain et al., 2020a). It is important for a CSR to perceive their coworker as trustworthy and respectful (Yang and Lau, 2019), even if it is AI (Bartneck et al., 2009).The current iteration of Pro-Pilot did not include any major anthropomorphic aspects to delineate the value of the support messages. Subsequent designs, however, can incorporate more anthropomorphic aspects (e.g., name, appearance, and tonality) to improve the emotional connectedness between the CSR and the AI-coworker (Hermann, 2022). At the same time, personalizing the conversations to users can make them more dynamic and human-like (Section 4.2.1). One way to do this is by learning from the CSR and other CSRs using long-term memory (LTM). Jo et al. found that AI agents that leverage LTM were perceived to be more personal and emotionally supportive (Jo et al., 2024). Beyond personalizing, LTM could also play a role in referring the CSR to real human stories retrieved from its memory. After all, many workers learn “on-the-job,” which is a form of social learning, or learning by observing those in one’s social group (Shteynberg and Apfelbaum, 2013). Future iterations could also be trained on human experiences that workers disclose on public social media (Saha et al., 2019; Das Swain et al., 2020b).. These practices are already common and unique to human coworker social networks (Section 6.3). AI-coworkers could act as a medium to share these experiences between CSRs and help them learn ER through each other.
Scaffolding Emotions in Task-Adjacent Moments: The lack of interventions for on-task cognitive reframing (Slovak et al., 2023) motivated us to design Pro-Pilot as an assistant that can intercept conversations with empathetic suggestions (Section 3). Instead, front-office work incorporates off-task ER methods. Some of these are infrastructural (e.g., training modules) and others are socio-cultural (e.g., internal forum for venting). We believe Pro-Pilot can not only co-exist with these off-task approaches, but it can also help bridge both approaches. For CSRs who regularly interact with the same clients, Pro-Pilot could learn from conversations to create a form of emotional briefing. This early information could act similarly to the usage of AI for planning, collaboration, and communication (Morrison et al., 2024). In turn, it can also be combined with AI for scheduling tasks and ensure workers can focus on important client-interactions when they have maximum resources available (Das Swain et al., 2023b) and also have opportunities to synchronize with their coworkers (Das Swain et al., 2019a). After the task, Pro-Pilot could also provide an emotional debriefing after an intense conversation to help the CSR recover in a way that resembles a coach. Software engineers are already appropriating AI for post-task recommendations (Cabrero-Daniel et al., 2024). What Pro-Pilot learns within the task can be extrapolated in between tasks to help CSRs smoothly ramp in and out of emotional labor.
7.2. Socio-Organizational Implications
Training, Evaluation & Safety. The current norm of desensitization to client incivility is partly because of the normative expectations that certain individuals are “built” for the job (e.g., “needs a thick skin”). Traits, however, do not develop overnight. Many participants saw Pro-Pilot as an early career training tool (Section 6.2). Learning modules could include Pro-Pilot’s simulation exercises to train ER (Slovak et al., 2023). A byproduct of many training is quantitative assessment. AI agents have already been proposed to assess worker mental health (Hungerbuehler et al., 2021). Workplaces are always looking to optimize their personnel through different algorithmic management methods but may not take the most responsible approaches (Das Swain et al., 2023a). Tools like Pro-Pilot do raise the risk of being appropriated to measure emotional labor. Kaur et al. has shown that some of the state-of-the-art methods to recognize worker emotion lack the necessary knowledge to accurately estimate a worker’s mental state (Kaur et al., 2022). Ironically, a CSR might end up performing more emotional labor to comply with Pro-Pilot’s expectations (Roemmich et al., 2023; Kawakami et al., 2023). Sometimes, a CSR should not tolerate client incivility (e.g., when a client uses obscene language and threatens harm). Organizations often have safeguards and protocols for these scenarios and Pro-Pilot needs to be integrated with such institutional knowledge. Using Pro-Pilot as a training tool is tempting but must be accompanied by appropriate protections and guidelines for use.
Mediating Social Relationships at Work. The messages generated by Pro-Pilot could be effective (Section 4.2.2), but it does not share the narrative style of humans (Section 4.2.2). On-task, a CSR may prefer Pro-Pilot because it relieves the burden of sharing context with a coworker to receive support (Section 6.2). From a utilitarian perspective, Pro-Pilot has advantages over human coworkers, but social relationships with coworkers offer more than just work-related support. The integration of Pro-Pilot into work might discourage CSRs from pro-social behaviors. Buçinca et al. have found evidence that AI assistance at work can erode social relationships (Buçinca et al., 2023). Especially given the anxiety around over-reliance on AI (Buçinca et al., 2021), one might be concerned about workplace social ties after empathetic AI coworkers become commonplace. One way forward, to avoid these concerns, is to augment AI coworkers by designing interventions that encourage human–human support, e.g., ”Consider reaching out to [your coworker], who went through a similar experience last week.” We urge future researchers to design them to not just mitigate negative social influences (e.g., incivility) but also promote positive social influences (e.g., peer support).
7.3. Limitations & Future Work
We create Pro-Pilot to unearth the potential of LLM–powered AI-coworkers as sources of emotional support in front-office work. Through our evaluations, we provide evidence on the effectiveness of AI-generated support messages and the usability of these insights in client incivility scenarios. All of our evaluations centered on the perspectives of real CSRs; however admittedly, transferring our findings directly to operational use needs more testing. Another important limitation of our study is that the effects of ER with Deep Acting, specifically Cognitive Change, only become apparent in the long term. Our results motivate field deployments and naturalistic testing for this purpose by demonstrating that Pro-Pilot has the feasibility and usability to anticipate improvement to long-term mental health at work. Researchers attempting these studies still need to overcome several non-trivial challenges. Future endeavors would need researchers to establish long-term partnerships with organizations that employ CSRs. Moreover, researchers need to consider the ethics and pragmatics of deploying such an AI-coworker intervention for critical roles such as those in front-office work. Even prior to field deployment, researchers can improve on our study on a few dimensions. The current iteration of emotional support messages is intentionally distinct from informational support. New iterations can explore and evaluate the value of combining these messages, i.e., conveying information with empathy. The user evaluation simulation study can be replicated at scale and control for order effects by recruiting enough people for between-subjects experiments. Our study design was tailored for incidents of client incivility, but future versions of Pro-Pilot need to be able to detect incivility, CSR’s affective response, and their goals to become truly dynamic.
8. Conclusion
Work as we know it is changing because of the advent of LLMs. AI applications are augmenting problem solving at a meteoric rate. However, front-office work involves more than procedural tasks. AI assistants need to support the emotional labor involved in work. Our study presents Pro-Pilot, an LLM-powered AI assistant to support the emotional labor of CSRs. We found that Pro-Pilot’s support messages were effective in expressing empathy appropriately (RQ1). We also found that CSRs could regulate their emotions thanks to Pro-Pilot’s Emo-Reframe function. These results open new doors for implementing holistic AI-coworkers and also raise important questions for socio-organizational development around this technology.
Appendix A Appendix
A.1. Generating Incidents with Uncivil Clients
A.1.1. Complaint categories
To enhance the realism of the simulation, we expanded the complaint categories using insights from previous customer complaint analyses (Cho et al., 2002). Furthermore, we refined these categories by incorporating a real-world dataset of publicly available client complaints from Twitter (Axelbrooke, 2017). These updated complaint categories were then used as prompts for GPT-4o, allowing for the creation of lifelike, multi-turn text interactions to simulate realistic customer service scenarios.
  1. (1)
    Service Quality: Issues related to human-to-human service interactions, such as delays, staff behavior, and communication errors.
  2. (2)
    Product Issues: Concerns regarding physical or functional aspects of a product or service, including defects, safety issues, and mismatches between expectation and reality.
  3. (3)
    Pricing and Charges: Financial discrepancies encountered before, during, or after a service, such as overcharging, undisclosed fees, or refund issues.
  4. (4)
    Policy: Grievances associated with company rules and guidelines, particularly when they are perceived as unfair or inflexible.
  5. (5)
    Resolution: The efficacy of the company’s actions in addressing and resolving complaints, focusing on customer satisfaction with the outcomes provided.
A.1.2. Behavioral context
Previous research examined how work engagement and challenge relate to focus, boredom, and routine tasks (Mark et al., 2014; Morshed et al., 2022). Based on these studies, we randomly assigned descriptions of the CSR’s mental state to two incidents.
  1. (1)
    Focused: “The conversation takes place about 2 hours into the work shift. The representative has already addressed a few customer complaints before the following incident.“
  2. (2)
    Stressed: “The conversation takes place in the second half of the work shift. The representative has been working longer hours over the past few days and has not been taking breaks.”
  3. (3)
    Bored: “The conversation takes place in the middle of the work shift. The representative has been spending minimal time on tasks and has been regularly checking their personal messages.”
A.1.3. Personality context
Personality affects how workers interpret and deal with situations (Donnellan and Robins, 2010). Participants were asked to recall a real coworker and choose one of three personality types. Descriptions of these traits were included in two incidents for added context.
  1. (1)
    Resilient: “They are organized and dependable. They tend to remain composed when facing challenges, but are prone to setting unrealistic expectations.”
  2. (2)
    Undercontrolled: “They are outgoing, competitive, and high energy. They tend to work on impulse, but are also prone to frustration.”
  3. (3)
    Overcontrolled: “They are detail-oriented and reliable but might appear distant. They tend to work carefully, but are prone to overthinking.”
A.2. User Evaluation: Survey Measurements
A.2.1. In-task Measurement
Throughout the study session, the participant was exposed to different insights from their AI coworker, Pro-pilot. To ensure participants have read these insights, they were required to answer a single-item question within each of these insight panels which we adapted from Samter et al.’s Effectiveness instrument and  Liu and Sundar’s semantic differential (Samter et al., 1987; Liu and Sundar, 2018). Fig. 4 shows how this was embedded into the task interface.
A.2.2. Post-task Measurement and Pre-task Measurement
After participants selected a scenario, participants were required to complete a pre-task survey to establish baseline attitudes. Upon concluding the conversation with each client, the participants were asked to reflect on the entire interaction and complete a post-task survey. The post-task survey contained additional questions, including Q4 (listed below), which was not present in the pre-task survey:
Client Interaction:
Q1.:
To what extent do you agree with the following statements about the client you conversed with (adapted from Spencer and Rupp(Spencer and Rupp, 2009)):
Strongly disagree $\cdot$ Disagree $\cdot$ Somewhat disagree $\cdot$ Neither agree nor disagree $\cdot$ Somewhat agree $\cdot$ Agree $\cdot$ Strongly agree
  1. (1)
    The client treated me in a polite manner.
  2. (2)
    The client treated me with dignity.
  3. (3)
    The client treated me with respect.
Cognitive Demands/Resources. We adapted both questions from Demerouti et al., which were rated on a 5-point Likert scale (Demerouti et al., 2001):
Very low $\cdot$ Low $\cdot$ Moderate $\cdot$ High $\cdot$ Very High
Q2a.:
In the context of your last conversation, how would you rate the demands on you?
Q2b.:
In the context of your last conversation, how would you rate the resources available to you?
Affect. We adapted two questions from  Betella and Verschure, scored on a semantic differential scale (Betella and Verschure, 2016):
Q3.:
How do you feel after the conversation?
a.:
Rate your level of pleasure
Negative———–————–————–————–————– Positive
b.:
Rate your level of energy
Calm———–————–————–————–————– Excited
Emotional Support. We adapted the Effectiveness and Supportiveness scales from  Liu and Sundar (Liu and Sundar, 2018). Each question was a 5-point semantic differential item; we list the extreme poles below:
Q4.:
How did you feel after reading the messages from Pro-pilot?
  • Effective / Ineffective
  • Helpful / Unhelpful
  • Beneficial / Not Beneficial
  • Adequate / Inadequate
  • Sensitive / Insensitive
  • Caring / Uncaring
  • Understanding / Not Understanding
  • Supportive / Unsupportive
A.3. Technical Evaluation: Robustness
Table A1. Summary of comparing the responses by Pro-Pilot and various LLMs (GPT-4, GPT-4o, Llama-3.1, and Mistral-7B in terms of $KW$-test (* $p$0.05, ** $p$¡0.01, *** $p$¡0.001). Only statistically significant results are reported.
[…]
Summary of results from comparison of reframes between Pro-Pilot and different cutting-edge LLMs. Categories PPilot GPT-4 GPT-4o Llama-3.1 Mistral-7B $H$-stat. Lexico-Semantics Syntax and Structure Verbosity 57.46 362.19 369.57 165.33 123.91 948.49*** Repeatability 0.20 0.45 0.45 0.39 0.34 635.98*** Readability 16.44 12.74 12.72 10.95 11.49 1931.01*** Linguistic Style and Semantics CDI 14.81 12.81 12.90 22.28 26.97 2592.60*** Empathy 0.91 0.87 0.87 0.77 0.88 375.29*** Adaptability 0.81 0.82 0.82 0.88 0.88 2933.78*** Psycholinguistics Affect Pos. Affect 0.043 0.052 0.053 0.018 0.019 1982.77*** Anger 0.020 0.005 0.005 0.008 0.006 872.04*** Sad 0.001 0.003 0.003 0.004 0.005 390.47*** Interpersonal Focus (Pronouns) 1st P. Sin. 0.005 0.016 0.015 0.028 0.029 1545.99*** 1st P. Plu. 0.002 0.034 0.035 0.004 0.002 2507.44*** 2nd P. 0.051 0.041 0.041 0.027 0.027 1119.48*** 3rd P. Plu. 0.038 0.013 0.011 0.009 0.007 993.46*** Impersonal Prn. 0.039 0.048 0.045 0.039 0.039 306.02***